OpenStudy (anonymous):

Derivative of y=x(x^2+3x)^1/2

OpenStudy (anonymous):

what did you try so far?

OpenStudy (anonymous):

Tried the chain rule but I messed it up somewhere. New to using it :(, I think we need to use chain rule more than once?

OpenStudy (anonymous):

can you write the general form of a derivative?

OpenStudy (anonymous):

P to the product

OpenStudy (anonymous):

like d/dx x^n = ?

OpenStudy (anonymous):

You have to use product rule

OpenStudy (anonymous):

\[fg = f'g+g'f\]

OpenStudy (anonymous):

Probably more than one way supposed to use chain rule though

OpenStudy (anonymous):

Yes, you have to use chain rule within the product rule

OpenStudy (welshfella):

its the product of 2 functions

OpenStudy (anonymous):

\[x \sqrt{x^2+3x}\] is this the problem?

OpenStudy (anonymous):

If so you have to use product rule and chain rule

OpenStudy (welshfella):

you use the chain rule to differentiate the radical part

OpenStudy (anonymous):

x(x^2+3x)^(1/2) yeah. Okay let me try hold on, thanks friends

OpenStudy (anonymous):

x*(1/2(x^2+3x))^(-1/2)*2x+3+(x^2+3x)^(1/2) Is what I got but it's wrong, anyone see my error?

OpenStudy (idku):

yes,

OpenStudy (idku):

\[x\cdot\frac{1}{2}(x^2+3x))^{-1/2}\cdot(2x+3)+(x^2+3x)^{1/2}\]

OpenStudy (idku):

maybe they want you to write it with positive exponents?

OpenStudy (anonymous):

Nope that shouldn't be the problem. We have the same answer right?